Let. (Matrix)

2             0             1

2             -2           2

0             4             1


(a) Find A2 and A3 , and verify that: A 3A 2 − 12A = −12i

holds, where I stands for the identity matrix.


(b) Find A−1 by multiplying the equation above on both sides by A−1

1
Expert's answer
2021-12-14T19:38:43-0500

(a) Let A = "\begin{bmatrix}
 2 & 0 &1\\
 2 & -2&2\\
0&4&1
\end{bmatrix}"

So A²

= "\begin{bmatrix}
 2 & 0 &1\\
 2 & -2&2\\
0&4&1
\end{bmatrix}\begin{bmatrix}
 2 & 0 &1\\
 2 & -2&2\\
0&4&1
\end{bmatrix}"


= "\begin{bmatrix}
 4+0+0 & 0+0+4 &2+0+1\\
 4-4+0 & 0+4+8&2-4+2\\
0+8+0&0-8+4&0+8+1
\end{bmatrix}" = "\begin{bmatrix}
 4 & 4 &3\\
 0 & 12&0\\
8&-4&9
\end{bmatrix}"

A³ = A²A

="\begin{bmatrix}
 4 & 4 &3\\
 0 & 12&0\\
8&-4&9
\end{bmatrix}" "\begin{bmatrix}
 2 & 0 &1\\
 2 & -2&2\\
0&4&1
\end{bmatrix}"

= "\begin{bmatrix}
 8+8+0 & 0 -8+12&4+8+3\\
 0+24+0 & 0-24+0&0+24+0\\
16-8+0&0+8+36&8-8+9
\end{bmatrix}"

= "\begin{bmatrix}
 16& 4&15\\
 2 4& -24&24\\
8&44&9
\end{bmatrix}"

Now A³ - A² =

"\begin{bmatrix}
 16& 4&15\\
 2 4& -24&24\\
8&44&9
\end{bmatrix}" "-\begin{bmatrix}
 4 & 4 &3\\
 0 & 12&0\\
8&-4&9
\end{bmatrix}"

= "\begin{bmatrix}
 12 & 0 &12\\
 24 & -36&24\\
0&48&0
\end{bmatrix}"


12A =

"12\begin{bmatrix}
 2 & 0 &1\\
 2 & -2&2\\
0&4&1
\end{bmatrix}" ="\begin{bmatrix}
 2 4& 0 &12\\
 24 & -24&24\\
0&48&12
\end{bmatrix}"

A³ - A² - 12A =

"\begin{bmatrix}
 12 & 0 &12\\
 24 & -36&24\\
0&48&0
\end{bmatrix}" "-\begin{bmatrix}
 2 4& 0 &12\\
 24 & -24&24\\
0&48&12
\end{bmatrix}"

= "\begin{bmatrix}
 -12& 0 &0\\
 0 & -12&0\\
0&0&-12
\end{bmatrix}"

= "-12\begin{bmatrix}
 1& 0 &0\\
 0 & 1&0\\
0&0&1
\end{bmatrix}= -12I"


Relation is verified.

b)

A³ - A² - 12A =-12I

Multiplying both sides by A-1

A³A-1- A²A-1-12AA-1=-12IA-1

=> A²I - AI - 12I = -12A-1

=> A² - A - 12I = -12A-1

=> -12A-1 = "\begin{bmatrix}
 4 & 4 &3\\
 0 & 12&0\\
8&-4&9
\end{bmatrix}" - "\begin{bmatrix}
 2 & 0 &1\\
 2 & -2&2\\
0&4&1
\end{bmatrix}" -"\begin{bmatrix}
 12 & 0 &0\\
 0 & 12&0\\
0&0&12
\end{bmatrix}"

= "\begin{bmatrix}
 4-2-12 & 4-0-0 &3-1-0\\
 0-2-0 & 12+2-12&0-2-0\\
8-0-0&-4-4-0&9-1-12
\end{bmatrix}" ="\begin{bmatrix}
 -10 & 4 &2\\
 -2 & 2&-2\\
8&-8&-4
\end{bmatrix}"

A-1 = "\frac{1}{6}\begin{bmatrix}
 5 & -2 &-1\\
 1 & -1&1\\
-4&4&2
\end{bmatrix}"